OpenStudy (anonymous):

Solve 2x^2 – 8x = –7.

hero (hero):

Hint: 2x^2 - 8x = -7 2(x^2 - 8x) = -7 x^2 - 8x = -3.5

OpenStudy (anonymous):

 x=2 cause u have to round it off.